Is the product of 11 and multiple of 101 is a number in which every digit is same?
Found 2 solutions by htmentor, Alan3354:
Answer by htmentor(1343)

The product of 11 and a multiple of 101 can be written as:
11*n*101 = 11n(100 + 1) where n is any integer
So the product is 1100n + 11n -> nn00 + nn = nnnn

Answer by Alan3354(69443)

Is the product of 11 and multiple of 101 is a number in which every digit is same?
-----------
It is if the (integral) multiple of 101 is less than 10.
For 10: 11*101*10 = 11110. Four 1's and a zero.